OpenStudy (anonymous):

in Bristol zoo there are 20 monkeys,10 penguins,8 lions,8 seals,6 tigers. what proportion of the animals are a monkeys b seals c tigers

OpenStudy (anonymous):

Given ; 20 monkeys 10 penguins 8 seals 8 lions 6 tigers Total = 20 + 10 + 8 + 8 + 6 =52 proportion of monkeys = 20/52 = 5/13 Proportion of seals = 8/52= 2/13 propryion of tigers =6/52= 3/26